OpenStudy (anonymous):

If 2a^2 = b^2 where a, b belong to integers, then 2 is a common divisor of a and b prove

OpenStudy (anonymous):

yikes this one again

OpenStudy (anonymous):

first of all it is not possible to have integer a and b with \[2a^2=b^2\] but that is ok, we can pretend here we go

OpenStudy (anonymous):

u cant really search for previous questions is sux

OpenStudy (anonymous):

first note that \[2a^2\] is even because it has a 2 in front

OpenStudy (anonymous):

then since \[2a^2=b^2\] that means \[b^2\] is even

OpenStudy (anonymous):

now if \[b^2\] is even , that means b is even because if you square an even number you get an even number, and if you square an odd number you get an odd number. so not only is b^2 even but also b must be even

OpenStudy (anonymous):

ok that was so easy idk why i didnt see it ty :)

OpenStudy (anonymous):

now since b is even, you know that not only is b^2 even, but it is divisible by 4, because the square of any even number is divisible by 4 (easy proof if you need it)

OpenStudy (anonymous):

well we are not quite done actually, because your question (although i don't actually see it) is to show that both b and a are even, so we have still to show that a is even

OpenStudy (anonymous):

we know that b is even, and that b^2 is divisible by 4 and we know \[2a^2=b^2\] which means \[2a^2\] is divisible by 4, which means \[a^2\] is even which means \[a\] is even as above. done

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

can u say that if 2|b, then there exist k such that 2k=b, so 2a^2 = (2k)^2 => a^2=2k^2, thus two is a common divisor of a and b?

OpenStudy (anonymous):

yw but please be advised it is not possible for \[2a^2=b^2\] for whole umber a and b

OpenStudy (anonymous):

@onceupon a time yes you are right, but you are assuming that b is even. you have to show that as well. i just said it in english. but the whole thing is stupid because you cannot have one whole number squared as twice another one squared. a picture will show you why

OpenStudy (anonymous):